OpenStudy (anonymous):

help solve for x 1/x=8/5x+3

OpenStudy (anonymous):

\[\frac{ 1 }{ x }=\frac{ 8 }{ 5x }+3\]

OpenStudy (anonymous):

Dunno if this is the most efficient way but i would subtract 8/5x from both sides... so you got (1/x)-(8/5x)=3 can you get them with common denominator?

OpenStudy (anonymous):

i got 0=8x^2+15x-5

OpenStudy (anonymous):

Doesnt look right... common denominator is 5x, you have to multiply top and bottom of first fraction to get the common denominator so (5/5x)-(8/5x)=3 -3/5x=3 now cross multiply

OpenStudy (anonymous):

where did 5/5x come from

OpenStudy (anonymous):

you want a common denominator so you can subtract 1/x - 8/5x common denominator is 5x and to get that in the first fraction you have to multiply top and bottom by 5 5/5x - 8/5x

OpenStudy (anonymous):

got it ! so how would i do -3/5x=3 to get x

OpenStudy (anonymous):

think of 3 as 3/1 and cross multiply -3/5x=3/1 -3=15x

OpenStudy (anonymous):

omg thank u so much bless ur soul

OpenStudy (anonymous):

You're welcome